A normal distribution has μ = 80 and σ = 10. What is the probability of randomly selecting between the mean and a score of 50


1
Expert's answer
2021-06-04T06:20:54-0400

"z_2=\frac{x_2-\mu}{\sigma}==\frac{80-80}{10}=0"


"z_1=\frac{x_1-\mu}{\sigma}=\frac{50-80}{10}=-3.00"


"P(z<0)=0.5"

"P(z<-3.00)=0.00135"


"P(50<x<80)=P(z<0)-P(z<-3.00)=0.5-0.00135=0.49865"